# Adaptive Massively Parallel Coloring in Sparse Graphs

### v1

@@ Abstract
Classic symmetry-breaking problems on graphs have gained a lot of attention in models of modern parallel computation. The Adaptive Massively Parallel Computation (AMPC) is a model that captures central challenges in data center computations. Chang et al. [PODC’2019] gave an extremely fast, constant time, algorithm for the $(\Delta+1)$-coloring problem, where $\Delta$ is the maximum degree of an input graph of $n$ nodes. The algorithm works in the most restrictive low-space setting, where each machine has $n^{\delta}$ local space for a constant $10<\delta<$ 1. The standard approaches for $(\Delta+1)$-coloring are ignorant about the graph topology in the following sense: They exploit the property that any partial coloring can be extended to a feasible $(\Delta+1)$-coloring of the whole graph. For most graphs, the chromatic number is much smaller than $\Delta+1$ and we would like to find colorings with fewer colors. However, as soon as we have fewer than $\Delta+1$ colors, completing partial colorings might not be possible. In this work, we study the vertex-coloring problem in sparse graphs parameterized by their arboricity $\alpha$, a standard measure for sparsity. We give deterministic algorithms that in constant, or almost constant, time give $\operatorname{\text{{\rm poly}}}\alpha$ and $O(\alpha)$-colorings, where $\alpha$ can be arbitrarily smaller than $\Delta$. A strong and standard approach to compute arboricity-dependent colorings is through the Nash-Williams forest decomposition, which gives rise to an (acyclic) orientation of the edges such that each node has a small outdegree. Our main technical contribution is giving efficient deterministic algorithms to compute these orientations and showing how to leverage them to find colorings in low-space AMPC. A key technical challenge is that the color of a node may depend on almost all of the other nodes in the graph and these dependencies cannot be stored on a single machine. Nevertheless, our novel and careful exploration technique yields the orientation, and the arboricity-dependent coloring, with a sublinear number of adaptive queries per node.
@@ 1 Introduction
In this paper, we study the graph coloring problem in sparse graphs in a modern model of parallel computation, the AMPC-Model [${}^{+}$19]. The Adaptive Massively Parallel Computation (AMPC) model of computation captures the challenges in modern platforms of processing massive data. We have $P$ machines with $S=O(n^{\delta})$ local space each, where $10<\delta<$ 1 and $n$ is the number of nodes in the input graph. In synchronous rounds, the machines can communicate through a distributed key-value storage. The machines have read access to the storage and, at the end of each round, the machines can write to the storage. A crucial limitation is that the local space limits the number of reads and writes per round. The AMPC model is a practically motivated extension of the standard MPC model, where the communication is restricted to all-to-all synchronous message-passing [KSV10].
Before delving into our results, let us explain the necessary background. The goal of the $k$-vertex-coloring problem is to assign a color out of $k$ colors to every node of a graph such that no neighboring nodes have the same color. This is a classic and central problem that has been intensively studied for decades. Vertex-coloring also has close connections to other classic problems such as the maximal independent set and scheduling.
In the distributed and parallel setting, the typical goal is to color the graph with $\Delta+1$ colors, where $\Delta$ corresponds to the maximum degree of the input graph. Indeed, any graph can be colored with $\Delta+1$ colors, and for cliques and odd cycles, this is tight. But many graphs can be colored with significantly fewer than $\Delta+1$ colors. For example, sparse graphs can have a high maximum degree but can often be colored with much fewer colors. Sparse graphs often occur in practice providing a demand for better coloring algorithms for sparse graphs.
@@
In this work, we study the vertex-coloring problem in graphs parameterized by arboricity, which is a standard measure of “everywhere” sparsity. The arboricity $\alpha$ corresponds to the minimum number of forests that the edges of a graph can be partitioned into [NW64]. It is well-known that any graph can be colored with $2\alpha$ colors and in some cases, this is tight. As a straightforward example, a tree clearly has arboricity $1$ and can be colored with $2$ colors, while the maximum degree $\Delta$ can be up to $n-1$. Hence, arboricity captures the number of colors needed to color sparse graphs.
@@
Finding arboricity-dependent colorings with $O(\alpha)$, $\operatorname{\text{{\rm poly}}}(\alpha)$, or $\exp(\alpha)$ colors has gained attention in various models of computation like LOCAL [BE10b, BE10a, Bar15, FHK16a, MT22, Mau21], MPC [GS19, FGG23, BCG20], dynamic algorithms [HNW20, CNR23, CR22, BCPS23], and streaming [BCG20]. A common approach is to find an acyclic orientation of the edges such that each node has an outdegree of $O(\alpha)$ or $\operatorname{\text{{\rm poly}}}\alpha$. Starting from sinks, the nodes can then be iteratively colored with outdegree + 1 colors. Typically, the outdegree gives a trade-off between the runtime and the number of required colors. This approach to coloring highlights a crucial difference between $(\Delta+1)$-coloring and arboricity-dependent coloring and pinpoints a key challenge. In the former case, any partial solution can be extended to a complete solution but this is not the case for the latter. In this light, the iterative coloring approach can be seen as carefully choosing partial solutions that can be extended to a complete solution.
@@
In recent work, the problem was studied in the context of dynamic algorithms [CNR23]. They showed that given such an orientation, there is an efficient Local Computation Algorithm (LCA) to find an arboricity-dependent coloring. They also showed how to maintain such an orientation efficiently. In the LCA model, each node can be queried for the $i$th entry in their adjacency list and their degree [RTVX11]. The goal is to implement an oracle, that returns the output at any node $v$, say, the color of $v$ in the coloring problem. The crux is that the outputs (per node) must yield a consistent and feasible solution over the whole graph.
@@
In our setting, the orientation is not given for free and one of our contributions is an LCA to find an acyclic orientation with a small outdegree. As pointed out by the authors in [CNR23], there is a caveat to designing such an LCA. While there is no formal proof, it seems likely the problem has a high volume, i.e., any LCA needs to do linear (in $n$) queries in the worst case. To circumvent this obstacle, we design an LCA that orients almost all of the edges. We show that this LCA is sufficient for designing efficient AMPC algorithms to compute acyclic low outdegree orientations orienting all edges. Lastly, we obtain various AMPC coloring algorithms that employ different tradeoffs between their runtime and the number of colors used. We next detail these three contributions in order, (1) the LCA for computing “weak orientations”, (2) AMPC algorithms for computing acyclic low outdegree orientations, and (3) AMPC coloring algorithms. We emphasize that all of our algorithms are deterministic.
@@
Contribution 1: An LCA for computing a weak orientation. A $\beta$-partition of a graph is a partition of its vertex set into layers such that every node has at most $\beta$ neighbors in higher (or the same) layers. The size of the partition is the number of layers. This partition naturally gives rise to the desired orientation required for efficient graph coloring algorithms by orienting edges from smaller to larger layers while orienting them arbitrarily within the layers. On graphs with arboricity at most $\alpha$, it is known that such partitions with $O(\log n)$ layers exist when $\beta\geq(2+\varepsilon)\alpha$ for any constant $\varepsilon>0$ [BE10b]. Our core building block and also the largest challenge of our approach is the design of an LCA algorithm that computes a weak version of the desired partition. In Section 2 we underline the technical challenges related to obtaining such a result. The following simplified version of the main lemma assumes that $\beta$ and $\alpha$ are constant and that $\beta\geq(2+\varepsilon)\alpha$ holds for any $\varepsilon>0$. We denote by $G[S]$ the subgraph induced by node set $S$ in graph $G$.
@@
For any constant $\delta>0$ there is deterministic LCA algorithm that uses at most $O(n^{\delta})$ queries per node on a graph $G$ with arboricity $\alpha$ and assigns each node a layer from ${}\mathbb{N}\cup\{\infty\}$ such that the following holds: normal-$\triangleright$ There exists a subset $S\subseteq V$ containing at least a $1-1/n^{O(\delta)}$ fraction of vertices such that the layering of $S$ forms a $\beta$-partition of $G[S]$ with $O(\log_{\beta}n)$ layers. normal-$\triangleleft$
Contribution 2: AMPC Algorithms for finding acyclic low outdegree orientations. Our second contribution is designing low-space AMPC algorithms to find acyclic orientations in graphs with any arboricity $\alpha$. Ultimately, our AMPC algorithm iteratively uses the above LCA to find a complete partition and the number of nodes that are not layered (after one application of the LCA) dictates the round complexity of the AMPC algorithm.
@@
Contribution 3: Coloring Sparse Graphs in AMPC. We design AMPC algorithms for coloring with different tradeoffs between their runtime and the number of colors used.
@@
    An $\alpha^{2+\varepsilon}$-coloring in $O(1/\varepsilon)$ rounds.
@@
All of our coloring algorithms heavily exploit that we first compute a small outdegree orientation using Theorem 1.2 with different parameters. While Theorem 1.3 has various tradeoffs between the runtime and the number of colors, its main strength lies in the following corollary.
@@
One of our easiest results of Theorem 1.3, once we’re given the small outdegree orientation, is the one coloring with quadratic number of colors. Here, we essentially simulate the LOCAL model algorithm by Linial that, very efficiently, computes a $\beta^{2}$-coloring on a graph with an outdegree $\beta$ and only uses local information. The fact that the algorithm works in graphs with a small outdegree (instead of small maximum degree as usually cited) was first observed in [BE10b]. Understanding when one can go below $o(\alpha^{2})$ colors with low volume and/or round complexity seems fundamentally hard and is an intriguing open question (see, e.g.,  [BE13, FHK16b]).
@@
We emphasize that, in contrast to most prior algorithms in the AMPC model and also most efficient LCA algorithms for large degree graphs, all our algorithms are deterministic. We note that deterministic algorithms seem fundamentally more constrained than randomized ones. As an example, in the deterministic setting, even on a tree, it seems out of reach to design a sublinear LCA that finds the nearest leaf of each node. In a closely related VOLUME model [^3] A crucial difference between LCA and VOLUME is that in VOLUME the queries must form a single connected component., it is known that $\Omega(n)$ deterministic queries are required to solve closely related problems on trees [RS20]. Furthermore, there is an exponential gap between the randomized and the deterministic complexities.
@@
A particular challenge in massively parallel computation models is to efficiently deal with nodes whose degree is significantly larger than the local space of each machine, especially for deterministic algorithms. As we desire to have efficient algorithms for all ranges of parameters, we design an algorithm using orthogonal techniques. We note that the following theorem holds in the strictly weaker [${}^{+}$19] low-space massively parallel computing model (MPC).
@@
In the case of $\alpha>n^{\delta}$, we can still compute the low outdegree partitions as claimed in Theorem 1.2. Actually, the partition comes with the property that each layer also has a small maximum degree, and hence, we can apply Theorem 1.5 (with $x=\alpha^{O(\varepsilon)}=n^{O(1)}$) to color each layer independently with $O(\alpha^{1+\varepsilon})$ colors in $O(\log_{x}n)=O(1)$ time, using a separate color space for each layer. This is one of the ways, how we use Theorem 1.5 in the proof of Theorem 1.3 in the case of huge arboricity.
@@ 1.1 Related Work
In the LCA model, there is an algorithm for the classic $(\Delta+1)$-coloring with $\operatorname{\text{{\rm poly}}}\Delta\cdot\log n$ queries [${}^{+}$19]. Given that the local space is larger than a small polynomial of $\Delta$, we can turn the LCA into a randomized AMPC algorithm that runs in constant time. The current deterministic state-of-the-art originates from the LOCAL model of distributed message-passing. Using the graph-exponentiation technique, we can simulate a known $\operatorname{\text{{\rm poly}}}\log\log n$round algorithm [CLP18] in $O(\log\log\log n)$ rounds. This can be further derandomized to match the same runtime deterministically and similar techniques apply to $(\textrm{degree}+1)$-coloring as well [CDP21, CCDPM24]. In AMPC, we can do even better. The aforementioned algorithms consist of two parts: Solving the problem on low-degree instances in $O(\log\log\log n)$ time and high-degree instances in $O(\log^{*}n)$ time. The former part can easily be simulated in $O(1)$ rounds in the low-space AMPC even though not explicitly stated in the papers. The runtime then reduces to $O(\log^{*}n)$.
@@
The coloring problem parameterized by the arboricity $\alpha$ has enjoyed attention in other related models of distributed and parallel computation as well. In the LOCAL model of distributed message-passing [Lin92], the edges of an input graph correspond to communication links along which the nodes communicate in synchronous rounds. Barenboim and Elkin gave deterministic $O(\log n)$ and $O(\alpha\log n)$ round algorithms for computing $O(\alpha^{2})$ and $(2+\varepsilon)\alpha$-colorings, respectively. It is also known that $\Omega(\log n)$ rounds are required [Lin92] for arboricity-dependent colorings. On the randomized side, it is known how to obtain an $O(\alpha\log\alpha)$ coloring in $O(\log n)$ rounds and $O(\alpha)$-coloring in $O(\log n\cdot\min\{\log\alpha,\log\log n\})$ rounds [GL17]. We note that there are also $\operatorname{\text{{\rm poly}}}\log n$ time algorithms for finding $(1+\varepsilon)\alpha$-outdegree orientations, but these orientations need not to be acyclic [HSV21].
@@
In the CONGESTED CLIQUE model, we have $n$ machines that communicate in an all-to-all fashion. Each machine corresponds to a node in the input graph and initially knows its incident edges. The CONGESTED CLIQUE model is very close to the non-adaptive MPC model with linear memory per machine, i.e., $\Omega(n)$ words per machine. There are deterministic algorithms for $O(\alpha)$-colorings in both of the models that run in $O(1)$ rounds [FGG23]. Finally, through the rake-and-compress decomposition on trees, one can also obtain an orientation with an outdegree at most $2$. In a recent work, it was shown how to obtain this decomposition in $O(1)$ rounds of AMPC [HKSS22]. For the special case of forests $(\alpha=1)$, [${}^{+}$23] give a $3$-coloring algorithm that runs in $O(\log\log n)$ rounds and uses optimal global space. This algorithm is conditionally optimal in the MPC model.

### v2

@@ Abstract
Classic symmetry-breaking problems on graphs have gained a lot of attention in models of modern parallel computation. The Adaptive Massively Parallel Computation (AMPC) is a model that captures the central challenges in data center computations. Chang et al. [PODC’2019] gave an extremely fast, constant time, algorithm for the $(\Delta+1)$-coloring problem, where $\Delta$ is the maximum degree of an input graph of $n$ nodes. The algorithm works in the most restrictive low-space setting, where each machine has $n^{\delta}$ local space for a constant $10<\delta<$ 1.
The standard approaches for $(\Delta+1)$-coloring are ignorant about the graph topology in the following sense: They exploit the property that any partial coloring can be extended to a feasible $(\Delta+1)$-coloring of the whole graph. For most graphs, the chromatic number is much smaller than $\Delta+1$ and we would like to find colorings with fewer colors. However, as soon as we have fewer than $\Delta+1$ colors, it might not be possible to complete partial colorings.
In this work, we study the vertex-coloring problem in sparse graphs parameterized by their arboricity $\alpha$, a standard measure for sparsity. We give deterministic algorithms that in constant, or almost constant, time give $\operatorname{\text{{\rm poly}}}\alpha$ and $O(\alpha)$-colorings, where $\alpha$ can be arbitrarily smaller than $\Delta$. A strong and standard approach to compute arboricity-dependent colorings is through the Nash-Williams forest decomposition, which gives rise to an (acyclic) orientation of the edges such that each node has a small out-degree.
Our main technical contribution is giving efficient deterministic algorithms to compute these orientations and showing how to leverage them to find colorings in low-space AMPC. A key technical challenge is that the color of a node may depend on almost all of the other nodes in the graph and these dependencies cannot be stored on a single machine. Nevertheless, our novel and careful exploration technique yields the orientation, and the arboricity-dependent coloring, with a sublinear number of adaptive queries per node.
@@ 1 Introduction
In this paper, we study the graph coloring problem in sparse graphs in a modern model of parallel computation, the AMPC-Model [BDE+19]. The Adaptive Massively Parallel Computation (AMPC) model of computation captures the challenges in modern platforms of processing massive data. We have $P$ machines with $S=O(n^{\delta})$ local space each, where $10<\delta<$ 1 and $n$ is the number of nodes in the input graph. In synchronous rounds, the machines can communicate through a distributed key-value storage. The machines have read access to the storage and, at the end of each round, the machines can write to the storage. A crucial limitation is that the local space limits the number of reads and writes per round. The AMPC model is a practically motivated extension of the standard MPC model, where communication is restricted to all-to-all synchronous message-passing [KSV10].
Before delving into our results, let us explain the necessary background. The goal of the $k$-vertex-coloring problem is to assign a color out of $k$ colors to every node of a graph such that no neighboring nodes have the same color. This is a classic and central problem that has been intensively studied for decades. Vertex-coloring also has close connections to other classic problems, such as the maximal independent set and scheduling.
In the distributed and parallel setting, the typical goal is to color the graph with $\Delta+1$ colors, where $\Delta$ corresponds to the maximum degree of the input graph. Indeed, any graph can be colored with $\Delta+1$ colors, and for cliques and odd cycles, this is tight. However, many graphs can be colored with significantly fewer than $\Delta+1$ colors. For example, sparse graphs can have a high maximum degree but can often be colored with much fewer colors. Sparse graphs often occur in practice, which is causing a demand for better coloring algorithms for sparse graphs.
@@
In this work, we study the vertex-coloring problem in graphs parameterized by arboricity, which is a standard measure of “everywhere” sparsity. The arboricity $\alpha$ corresponds to the minimum number of forests that the edges of a graph can be partitioned into [NW64]. It is well known that any graph can be colored with $2\alpha$ colors, and in some cases this is tight. As a straightforward example, a tree clearly has arboricity $1$ and can be colored with $2$ colors, while the maximum degree $\Delta$ can be up to $n-1$. Hence, arboricity captures the number of colors needed to color sparse graphs.
@@
Finding arboricity-dependent colorings with $O(\alpha)$, $\operatorname{\text{{\rm poly}}}(\alpha)$, or $\exp(\alpha)$ colors has gained attention in various models of computation like LOCAL [BE10b, BE10a, Bar15, FHK16a, MT22, Mau21], MPC [GS19, FGG23, BCG20], dynamic algorithms [HNW20, CNR23, CR22, BCPS23], and streaming [BCG20]. A common approach is to find an acyclic orientation of the edges such that each node has an out-degree of $O(\alpha)$ or $\operatorname{\text{{\rm poly}}}\alpha$. Starting from sinks, the nodes can then be iteratively colored with out-degree + 1 colors. Typically, the out-degree gives a trade-off between the runtime and the number of required colors. This approach to coloring highlights a crucial difference between $(\Delta+1)$-coloring and arboricity-dependent coloring and pinpoints a key challenge. In the former case, any partial solution can be extended to a complete solution, but this is not the case for the latter. In this light, the iterative coloring approach can be seen as carefully choosing partial solutions that can be extended to a complete solution.
@@
In recent work, the problem has been studied in the context of dynamic algorithms [CNR23]. They showed that given such an orientation, there is an efficient Local Computation Algorithm (LCA) to find an arboricity-dependent coloring. They also showed how to maintain such an orientation efficiently. In the LCA model, each node can be queried for the $i$th entry in their adjacency list and their degree [RTVX11]. The goal is to implement an oracle that returns the output at any node $v$, say, the color of $v$ in the coloring problem. The crux is that the outputs (per node) must yield a consistent and feasible solution throughout the graph.
@@
In our setting, the orientation is not given for free and one of our contributions is an LCA to find an acyclic orientation with a small out-degree. As pointed out by the authors in [CNR23], there is a caveat to designing such an LCA. Although there is no formal proof, it seems likely that the problem has a high volume, that is, any LCA needs to do linear (in $n$) queries in the worst case. To circumvent this obstacle, we design an LCA that orients almost all of the edges. We show that this LCA is sufficient for designing efficient AMPC algorithms to compute acyclic low out-degree orientations orienting all edges. Lastly, we obtain various AMPC coloring algorithms that employ different trade-offs between their runtime and the number of colors used. We next detail these three contributions in order, (1) the LCA for computing “weak orientations”, (2) AMPC algorithms for computing acyclic low out-degree orientations, and (3) AMPC coloring algorithms. We emphasize that all our algorithms are deterministic.
@@
Contribution 1: An LCA for computing a weak orientation. A $\beta$-partition of a graph is a partition of its vertex set into layers such that every node has at most $\beta$ neighbors in higher (or the same) layers. The size of the partition is the number of layers. This partition naturally gives rise to the desired orientation required for efficient graph coloring algorithms by orienting edges from smaller to larger layers while orienting them arbitrarily within the layers. In graphs with arboricity at most $\alpha$, it is known that such partitions with $O(\log n)$ layers exist when $\beta\geq(2+\varepsilon)\alpha$ for any constant $\varepsilon>0$ [BE10b]. Our core building block and also the largest challenge of our approach is the design of an LCA algorithm that computes a weak version of the desired partition. In Section 2 we underline the technical challenges associated with obtaining such a result. The following simplified version of the main lemma assumes that $\beta$ and $\alpha$ are constant and that $\beta\geq(2+\varepsilon)\alpha$ holds for any $\varepsilon>0$. We denote by $G[S]$ the subgraph induced by the node set $S$ in graph $G$.
@@
For any constant $\delta>0$ there is deterministic LCA algorithm that uses at most $O(n^{\delta})$ queries per node on a graph $G$ with arboricity $\alpha$ and assigns each node a layer from ${}\mathbb{N}\cup\{\infty\}$ such that the following holds:
$\triangleright$ There exists a subset $S\subseteq V$ containing at least a $1-1/n^{O(\delta)}$ fraction of vertices such that the layering of $S$ forms a $\beta$-partition of $G[S]$ with $O(\log_{\beta}n)$ layers. $\triangleleft$
Contribution 2: AMPC algorithms for finding acyclic low out-degree orientations. Our second contribution is designing low-space AMPC algorithms to find acyclic orientations in graphs with any arboricity $\alpha$. Ultimately, our AMPC algorithm iteratively uses the above LCA to find a complete partition and the number of nodes that are not layered (after one application of the LCA) dictates the round complexity of the AMPC algorithm.
@@
Contribution 3: Coloring sparse graphs in AMPC. We design AMPC algorithms for coloring with different trade-offs between their runtime and the number of colors used.
@@
    An $O(\alpha^{2+\varepsilon})$-coloring in $O(1/\varepsilon)$ rounds.
@@
All of our coloring algorithms heavily exploit that we first compute a small out-degree orientation using Theorem 1.2 with different parameters. Although Theorem 1.3 has various trade-offs between the runtime and the number of colors, its main strength lies in the following corollary.
@@
One of our easiest results of Theorem 1.3, once we’re given the small out-degree orientation, is the one coloring with quadratic number of colors. Here, we essentially simulate the LOCAL model algorithm by Linial that, very efficiently, computes a $\beta^{2}$-coloring on a graph with an out-degree $\beta$ and only uses local information. The fact that the algorithm works in graphs with a small out-degree (instead of small maximum degree as usually cited) was first observed in [BE10b]. Understanding when one can go below $o(\alpha^{2})$ colors with low volume and/or round complexity seems fundamentally hard and is an intriguing open question (see, e.g.,  [BE13, FHK16b]).
@@
We emphasize that, in contrast to most prior algorithms in the AMPC model and also the most efficient LCA algorithms for large degree graphs, all our algorithms are deterministic. We note that deterministic algorithms seem fundamentally more constrained than randomized ones. As an example, in the deterministic setting, even on a tree, it seems out of reach to design a sublinear LCA that finds the nearest leaf of each node. In a closely related VOLUME model [^3] A crucial difference between LCA and VOLUME is that in VOLUME the queries must form a single connected component., it is known that $\Omega(n)$ deterministic queries are required to solve closely related problems on trees [RS20]. Furthermore, there is an exponential gap between the randomized and the deterministic complexities.
@@
A particular challenge in massively parallel computation models is to efficiently deal with nodes whose degree is significantly larger than the local space of each machine, especially for deterministic algorithms. As we desire to have efficient algorithms for all ranges of parameters, we design an algorithm using orthogonal techniques. We note that the following theorem holds in the strictly weaker [BDE+19] low-space massively parallel computing model (MPC).
@@
In the case of $\alpha>n^{\delta}$, we can still compute the low out-degree partitions as claimed in Theorem 1.2. Actually, the partition comes with the property that each layer also has a small maximum degree, and hence, we can apply Theorem 1.5 (with $x=\alpha^{O(\varepsilon)}=n^{O(1)}$) to color each layer independently with $O(\alpha^{1+\varepsilon})$ colors in $O(\log_{x}n)=O(1)$ time, using a separate color space for each layer. This is one of the ways, how we use Theorem 1.5 in the proof of Theorem 1.3 in the case of huge arboricity.
@@ 1.1 Related Work
In the LCA model, there is an algorithm for the classic $(\Delta+1)$-coloring problem with $\operatorname{\text{{\rm poly}}}\Delta\cdot\log n$ queries [CFG+19]. Given that the local space is larger than a small polynomial of $\Delta$, we can turn the LCA into a randomized AMPC algorithm that runs in constant time. The current deterministic state-of-the-art originates from the LOCAL model of distributed message-passing [Lin92], where the edges of an input graph correspond to communication links along which the nodes communicate in synchronous rounds. Using the graph-exponentiation technique, we can simulate a known $\operatorname{\text{{\rm poly}}}\log\log n$round algorithm [CLP18] in $O(\log\log\log n)$ rounds. This can be further derandomized to match the same runtime deterministically and similar techniques apply to $(\textrm{degree}+1)$-coloring as well [CDP21, CCDPM24]. In AMPC, we can do even better. The aforementioned algorithms consist of two parts: Solving the problem on low-degree instances in $O(\log\log\log n)$ time and high-degree instances in $O(\log^{*}n)$ time. The former part can easily be simulated in $O(1)$ rounds in the low-space AMPC even though not explicitly stated in the papers. The runtime then reduces to $O(\log^{*}n)$.
@@
The coloring problem parameterized by the arboricity $\alpha$ has enjoyed attention in other related models of distributed and parallel computation as well. In the LOCAL model, Barenboim and Elkin gave deterministic $O(\log n)$ and $O(\alpha\log n)$ round algorithms for computing $O(\alpha^{2})$ and $(2+\varepsilon)\alpha$-colorings, respectively. It is also known that $\Omega(\log n)$ rounds are required [Lin92] for arboricity-dependent colorings. On the randomized side, it is known how to obtain an $O(\alpha\log\alpha)$ coloring in $O(\log n)$ rounds and $O(\alpha)$-coloring in $O(\log n\cdot\min\{\log\alpha,\log\log n\})$ rounds [GL17]. We note that there are also $\operatorname{\text{{\rm poly}}}\log n$ time algorithms for finding $(1+\varepsilon)\alpha$-out-degree orientations, but these orientations need not be acyclic [HSV21].
@@
In the CONGESTED CLIQUE model, we have $n$ machines that communicate in an all-to-all fashion. Each machine corresponds to a node in the input graph and initially knows its incident edges. The CONGESTED CLIQUE model is very close to the non-adaptive MPC model with linear memory per machine, i.e., $\Omega(n)$ words per machine. There are deterministic algorithms for $O(\alpha)$-colorings in both of the models that run in $O(1)$ rounds [FGG23]. Finally, through the rake-and-compress decomposition on trees, one can also obtain an orientation with an out-degree at most $2$. In a recent work, it was shown how to obtain this decomposition in $O(1)$ rounds of AMPC [HKSS22]. For the special case of forests $(\alpha=1)$, [GLM+23] give a $3$-coloring algorithm that runs in $O(\log\log n)$ rounds and uses optimal global space. This algorithm is conditionally optimal in the MPC model.